The following data were collected on the height (inches) and weight (pounds) of 5 students.

Height

72

70

62

65

67

Weight

180

172

125

132

145

a. Develop a regression model to predict weight based on height.

b. What percent of the total variation in weight has been explained by height?

c. If a student is 69 inches tall, what would you estimate the weight to be?

Please use the Excel Solver to solve the above exercise question
Answer #1

a.

Using excel for performing regression with weight as the dependent variable and height as independent variables.

In excel ,go to data tab-->data analysis and select regression

In the regression dialog select the weight values as Y range and height values as X range and click OK.

We get the regression output with coefficients of height and intercept .

The regression equation is ,Weight =Intercept+Coefficient of height *height

Weight=-251.758+5.990*height

b.From the regression output,

Coefficient of determination =r^2 =0.9570

The percentage of total variation in Weight that can be explained by height is r^2=0.9570 =95.70 %

So,95.70 % is explained by the regression line.

c.If height =69 inches,

We have, Weight=-251.758+5.990*height=Weight=-251.758+5.990*69=161.58 =162 (rounding to nearest integer)